OpenStudy (anonymous):

solve for x: (√x-1)-(√x-6)= 1

OpenStudy (raden):

√(x-1) - √(x-6)= 1 √(x-1) = 1 + √(x-6) u have to squaring to both sides, what u get ...

OpenStudy (anonymous):

i got x=10

OpenStudy (raden):

yea, i got x=10 too

OpenStudy (anonymous):

k thanks

OpenStudy (raden):

yw